OpenStudy (anonymous):

f(x) = x^2 + x - 30 what are the zeroes of the function? I'm curious as to how to initially set this up so I can find the zeros. Anyone willing to guide me?

OpenStudy (softballgirl372015):

You have to factor this function. This factors to (x+6)(x-5)

OpenStudy (anonymous):

How do you factor it to get to that?

OpenStudy (softballgirl372015):

Have you learned the quadratic formula yet?

OpenStudy (anonymous):

I've learned a lot. Also forgotten a lot x) I'll go look it up.

OpenStudy (softballgirl372015):

If you don't know how to factor, you can also use the quadratic formula to find the zeros of the function. its \[-\frac{ b+ or -\sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (softballgirl372015):

In this problem, a=1 b=1 and c=-30

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

oh woops you already posted it lol

OpenStudy (anonymous):

okay so lemme think for a moment

OpenStudy (anonymous):

are a and b 1 by default if their value is unknown?

OpenStudy (anonymous):

@abb0t @AriPotta @AlexandervonHumboldt2 @acxbox22 @AndrewTheCookie @bahrom7893 @bohotness

OpenStudy (alexandervonhumboldt2):

so find the values of x in the equation using quadratic formula. then factor

OpenStudy (anonymous):

are a and b 1 by default if their value is unknown?

OpenStudy (anonymous):

That was what I was wondering

OpenStudy (softballgirl372015):

their values are known as 1 because 1 times any number is that number. The 1 is unnecessary to include in the coefficient because it is implied.

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

so then I say 1 +or - (sqrt 1^2 - (1*30)) / (2*1)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Softballgirl372015

OpenStudy (anonymous):

@Callisto @camerondoherty @cupcakerain @Chop-GTA5 @clamin @ChasingMaggie @DanJS @dtan5457 @dmndlife24 @DiamondBear3299 @FuyukoHarayo @falloutboi @Fireworktryingtopass! @fanny8

OpenStudy (anonymous):

@gottennis121 @GhostGamerYT @gottalottablue @greenman44 @gabgurl @girlover @hotgurl72 @hysusonic @HeyWassup345 @hayleynicole123 @heartbreaker_girl

OpenStudy (anonymous):

@iamabarbiegirl @ICECREAMyummy @IloveHW @Iloveyou;)boy @iPwnBunnies @iambatman @jim_thompson5910 @Jesstho.-. @jordanloveangel @Jaynator495 @jagr2713 @Jamielynne